# DynFocus: Dynamic Cooperative Network Empowers LLMs with Video Understanding

### v1

@@ Abstract
The challenge in LLM-based video understanding lies in preserving visual and semantic information in long videos while maintaining a memory-affordable token count. However, redundancy and correspondence in videos have hindered the performance potential of existing methods. Through statistical learning on current datasets, we observe that redundancy occurs in both repeated and answer-irrelevant frames, and the corresponding frames vary with different questions. This suggests the possibility of adopting dynamic encoding to balance detailed video information preservation with token budget reduction. To this end, we propose a dynamic cooperative network, DynFocus, for memory-efficient video encoding in this paper. Specifically, i) a Dynamic Event Prototype Estimation (DPE) module to dynamically select meaningful frames for question answering; (ii) a Compact Cooperative Encoding (CCE) module that encodes meaningful frames with detailed visual appearance and the remaining frames with sketchy perception separately. We evaluate our method on five publicly available benchmarks, and experimental results consistently demonstrate that our method achieves competitive performance.
((a))
((b))
@@ 1 Introduction
Figure 2: Competitive even SOTA performance on various video understanding tasks.
@@
  • Experimentally, we evaluate our DynFocus on two publicly mainstream short video benchmarks, three long video benchmarks, and one diagnosis benchmark on video hallucination as shown in Fig. 2.
@@ 3.1 Visual and Text Encoder
Given a $T$-frame video, we extract the frame-wise features $\mathbf{V}=\{\mathbf{v}_{t}\}_{t=1}^{T}$, using a pre-trained visual encoder, where $\mathbf{v}_{t}\in\mathbb{R}^{N\times d}$ denotes the feature of the $t$-th frame. Here, $N$ is the number of image patches and $d$ is the feature dimension. For the text encoder, user instruction is fed to the pre-trained text encoder to generate the text features $\mathbf{Q}=\{\mathbf{q}_{r}\}_{r=1}^{R}$, where $\mathbf{q}_{r}\in\mathbb{R}^{d^{\prime}}$ denotes the feature embedding of each word in user instruction. $d^{\prime}$ is the feature dimension and $R$ is the total number of words.
@@ 3.2 Dynamic Event Prototype Estimation
For clustering, the local density $\rho_{t}$ measures the mean distance to $K$ nearest neighbors of the $t$-th frame, and is given by,
$\rho_{t}=\mathrm{exp}\left(-\frac{1}{K}\sum_{{t^{\prime}}\in\mathcal{N}({t})}\frac{1}{P}\|\mathbf{f}_{t}-\mathbf{f}_{t^{\prime}}\|_{F}^{2}\right),$ (1)
where ${{t^{\prime}}\in\mathcal{N}({t})}$ denotes that $\mathbf{f}_{t^{\prime}}$ is in the neighborhood of $\mathbf{f}_{t}$. $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm, and it is also used to perform $K$ nearest neighbor search for each $\mathbf{f}_{t}$.
@@
where $\mathrm{Max}(\mathbf{m}_{l})\in\mathbb{R}^{d}$ denotes the row feature with maximum feature norm ($L_{2}$) across $P$ rows. $\mathrm{Avg}(\mathbf{m}_{l})\in\mathbb{R}^{d}$ denotes the averaged row feature across $P$ rows. |$\cdot$|$\cdot$ | | ⋅ denotes concatenation. After collecting all scores into a score vector $\mathbf{s}=\{\mathbf{s}_{l}\}_{l=1}^{L}$, we perform min-max normalization to normalize score values to be within $[0,1]$. Finally, we sort scores in the decreasing order and take the Top-$K$ event prototypes from $\mathbf{M}$ with high scores as filtered event prototypes, denoted by $\mathbf{H}=\{\mathbf{h}_{k}\}_{k=1}^{K}$. We also retrieve indices of filtered event prototypes in the $T$-frame video, obtaining the binary index mask $\mathbf{b}=\{b_{t}\}_{t=1}^{T}$. Notably, $b_{t}=1$ indicates important frame, while $b_{t}=0$ signifies redundant frame.
Training $\mathcal{U}(\cdot)$. Note that the Top-$K$ operation is not differentiable and thus stops the gradient propagation from LLM to update our score net $\mathcal{U}(\cdot)$. This limitation restricts our $\mathcal{U}(\cdot)$ to dynamically estimate the redundancy and flexibly capture correspondence without auxiliary loss supervision. To address this issue, we transform the Top-$K$ operation into solving a linear programming problem to make our network end-to-end trainable, and it can be directly updated by the autoregressive loss between language response and ground-truth answer within LLM. Please refer to the supplementary material for details.
@@ 3.3 Compact Cooperative Encoding
We find that semantic abstraction of $\mathbf{v}_{t}$ can be achieved when spatial clustering is performed multiple times. For example, concepts such as “person” and “dog” are progressively formed from the low-level attribute or color information. We thus use multiple clustering layers to capture more abundant visual details. The output spatial object prototypes of layer $j$ is fed to layer $j+1$ for abstraction. Collecting all layer outputs results in our multi-grained spatial object prototypes $\mathbf{G}_{t}=\{\mathbf{Z}^{(j)}_{t}\}_{j=1}^{J}$. Here, $J$ is the total number of clustering layers.
@@
Given embeddings $\mathbf{U}_{t,b_{t}=1}$ and $\mathbf{U}_{t,b_{t}=0}$ from Cones and Rods, respectively, we combine them to obtain the dynamic embedding of the $t$-th frame,
@@ 4.1 Experimental Setup
Implementation Details. We use the pre-trained ViT-G/14 from EVA-CLIP [12] as the visual encoder to extract the features of each frame in video, and it can be further changed to other clip-based video encoders. We use pre-trained Qformer weight from InstructBLIP [9] as the textual encoder. Besides, we adopt the Vicuna-7B model [7] as our foundational LLM. Our model is trained using 8 $\times$ NVIDIA A100 80G GPUs.
@@
[…]
Table 2: Performance comparisons on VCG-Bench. $\dagger$ represents the version that first fine-tuned on all the dataset, and further post-tuned on VideoChatGPT-100K [37] with a smaller learning rate.
@@
[…]
@@
VideoChat [25] 16 fr 33.0 32.0 2.31 27.0 32.1 27.6 5.01 24.3 28.6 29.2 3.66
Video-ChatGPT [ACL 24] [37] 100 fr 26.9 24.0 2.31 40.3 42.0 29.9 5.48 25.1 31.1 31.3 3.90
Video-LLaMA2 [6] 16 fr 54.5 41.5 2.34 39.4 33.5 35.4 5.22 18.5 25.7 35.5 3.78
[…]
@@ 4.2 Evaluation on Short Video Understanding
Video-LLaVA [EMNLP 24] [30] 8 fr 7B 45.3 46.1 38.0 40.7 36.2 38.1 39.9 41.6
ST-LLM [ECCV 24] [34] 16 fr 7B 45.7 48.4 36.8 41.4 31.3 36.9 37.9 42.3
[…]
@@
Table 5: Performance Comparisons on LV-Bench. Input shows the number of frames each model can process.
Method Size Input ER EU KIR TG Rea Sum Overall
Short Video MLLMs
TimeChat [CVPR 24] [44] 7B 96 fr 21.9 21.7 25.9 22.7 25.0 24.1 22.3
PLLaVA [61] 34B 16 fr 25.0 24.9 26.2 21.4 30.0 25.9 26.1
LLaVA-NeXT [65] 34B 32 fr 30.1 31.2 34.1 31.4 35.0 27.6 32.2
GPT-4o [41] - 10 fr 26.5 23.7 28.3 21.4 28.0 32.8 27.0
Long Video MLLMs
MovieChat [CVPR 24] [49] 7B $\sim$10k fr 21.3 23.1 25.9 22.3 24.0 17.2 22.5
LLaMA-VID [ECCV 24] [28] 13B $\sim$10k fr 25.4 21.7 23.4 26.4 26.5 17.2 23.9
LWM [33] 7B $\sim$4k fr 24.7 24.8 26.5 28.6 30.5 22.4 25.5
Gemini 1.5 Pro [43] - 3.6k fr 32.1 30.9 39.3 31.8 27.0 32.8 33.1
[…]
@@ 4.3 Evaluation on Long Video Understanding
To demonstrate the advantage of our dynamic cooperative setting, we conduct experiment on three newly released long-term video benchmark. The detailed description for each benchmark are elaborated in supplementary material. MLVU-Bench performance. The evaluation task of MLVU can be categorized into three types: (1) holistic LVU, which requires to make use of global perspectives from the entire video; (2) single-detail LVU, which needs to pinpoint one critical details in a haystack; (3) multi-detail LVU, which calls for the joint utilization of multiple detailed plots within the long video to collaborately infer the answer.
The performance of individual task and the average performance of multi-choice task (M-Avg, within 0-100%) and generation task (G-Avg, within 0.0-10.0) are both reported in Table 3. We have following observations: (1) our DynFocus surpasses all the open-sourced video MLLMs with a clear-cut performance gain on M-Avg and G-Avg, and it nearly consistently ranks top-2 position on individual tasks. (2) For TR, AR, and VS tasks that require an thorough understanding of entire video, our method achieves the best. We attribute this to our dynamic cooperative network’s ability to balance intricate spatial details with broader temporal perception without introducing external visual nuisance. (3) Most approaches find AO and AC tasks challenging due to their sensitivity to the temporal clues, which requires recalling multiple nuanced details from lengthy videos. Although not being further fine-tuned on long-term video dataset like MovieChat, our model still performs competitively. (4) However, our method struggles with ER task that needs ego-based perspectives, likely due to the requirement for ego-centric dataset like EgoQA [10] in VideoChat2.
LV-Bench Performance. We assess six core capabilities of our model on LV-Bench: Temporal Grounding (TG), Summarization (Sum), Reasoning (Rea), Entity Recognition (ER), Event Understanding (EU), and Key Information Retrieval (KIR). The average duration of each video exceeds 1 hour. Following [53], we select several publicly evaluated methods as baselines, with results shown in Table 5. Interestingly, some methods that excel on short videos perform almost randomly in answer selection. Remarkably, our DynFocus achieves the best of 31.5% among all the open-sourced 7B models, even outperforming PLLaVA [26] with 34B parameters.
((a))
((b))
@@
VideoMME Benchmark Performance. VideoMME benchmark spans across three kinds of durations, VideoMME-S ($\sim$1.3min), VideoMME-M ($\sim$8.5min), and VideoMME-L ($\sim$0.7h). Most videos include both subtitles and audios, which helps us investigate the performance gain from additional information sources. Figure 4 compares our results with other representative video MLLMs. Notably, our DynFocus consistently achieves impressive advantage across different lengths of video with subtitles and without subtitles. Specifically, it exhibits the overwhelming advantage over SOTA ST-LLM and VideoChat2. Remarkably, the version of DynFocus without subtitles reaches an overall accuracy of 44.1%, still surpassing ST-LLM with subtitles by 1.8%.

### v2

@@ Abstract
The challenge in LLM-based video understanding lies in preserving visual and semantic information in long videos while maintaining a memory-affordable token count. However, redundancy and correspondence in videos have hindered the performance potential of existing methods. Through statistical learning on current datasets, we observe that redundancy occurs in both repeated and answer-irrelevant frames, and the corresponding frames vary with different questions. This suggests the possibility of adopting dynamic encoding to balance detailed video information preservation with token budget reduction. To this end, we propose a dynamic cooperative network, DynFocus, for memory-efficient video encoding in this paper. Specifically, i) a Dynamic Event Prototype Estimation (DPE) module to dynamically select meaningful frames for question answering; (ii) a Compact Cooperative Encoding (CCE) module that encodes meaningful frames with detailed visual appearance and the remaining frames with sketchy perception separately. We evaluate our method on five publicly available benchmarks, and experimental results consistently demonstrate that our method achieves competitive performance. Code is available at https://github.com/Simon98-AI/DynFocus
@@ 1 Introduction
  • Experimentally, we achieve the competitive even SOTA performance on two publicly mainstream short video benchmarks, three long video benchmarks, and one diagnosis benchmark on video hallucination.
@@ 3.1 Visual and Text Encoder
Given a $T$-frame video, we extract the frame-wise features $\mathbf{V}=\{\mathbf{v}_{t}\}_{t=1}^{T}$, using a pre-trained visual encoder, where $\mathbf{v}_{t}\in\mathbb{R}^{N\times d}$ denotes the feature of the $t$-th frame. Here, $N$ is the number of image patches and $d$ is the feature dimension. For the text encoder, user instruction is fed to the pre-trained text encoder to generate the text features $\mathbf{Q}=\{\mathbf{q}_{r}\}_{r=1}^{R}$, where $\mathbf{q}_{r}\in\mathbb{R}^{d^{\prime}}$ denotes the feature embedding of each token in user instruction. $d^{\prime}$ is the feature dimension and $R$ is the total number of tokens.
@@ 3.2 Dynamic Event Prototype Estimation
For clustering, the local density $\rho_{t}$ measures the mean distance to $C$ nearest neighbors of the $t$-th frame, and is given by,
$\rho_{t}=\mathrm{exp}\left(-\frac{1}{C}\sum_{{t^{\prime}}\in\mathcal{N}({t})}\frac{1}{P}\|\mathbf{f}_{t}-\mathbf{f}_{t^{\prime}}\|_{F}^{2}\right),$ (1)
where ${{t^{\prime}}\in\mathcal{N}({t})}$ denotes that $\mathbf{f}_{t^{\prime}}$ is in the neighborhood of $\mathbf{f}_{t}$. $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm, and it is also used to perform $C$ nearest neighbor search for each $\mathbf{f}_{t}$.
@@
where $\mathrm{Max}(\mathbf{m}_{l})\in\mathbb{R}^{d}$ denotes the row feature with maximum feature norm ($L_{2}$) across $P$ rows. $\mathrm{Avg}(\mathbf{m}_{l})\in\mathbb{R}^{d}$ denotes the averaged row feature across $P$ rows. |$\cdot$|$\cdot$ | | ⋅ denotes concatenation. After collecting all scores into a score vector $\mathbf{s}=\{\mathbf{s}_{l}\}_{l=1}^{L}$, we perform min-max normalization to normalize score values to be within $[0,1]$. Finally, we sort scores in the decreasing order and take the Top-$K$ event prototypes from $\mathbf{M}$ with high scores, indicated by the index vector $\mathbf{p}$, as filtered event prototypes. The filtered event prototypes and the index vector are denoted as $\mathbf{H}=\{\mathbf{h}_{k}\}_{k=1}^{K}$ and $\mathbf{p}=\mathrm{topk}(\mathbf{s})\in\mathbb{N}^{K}$, respectively.
We also retrieve indices of filtered event prototypes in the $T$-frame video, obtaining the binary index mask $\mathbf{b}=\{b_{t}\}_{t=1}^{T}$. Notably, $b_{t}=1$ indicates important frame, while $b_{t}=0$ signifies redundant frame.
Training $\mathcal{U}(\cdot)$. Note that the Top-$K$ operation is not differentiable and thus stops the gradient propagation from LLM to update our score net $\mathcal{U}(\cdot)$. This limitation restricts our $\mathcal{U}(\cdot)$ to dynamically estimate the redundancy and flexibly capture correspondence without auxiliary loss supervision. To address this issue, we transform the Top-$K$ operation into solving a linear programming problem to make our network end-to-end trainable. Specifically, we convert the index vector $\mathbf{p}=\left[p_{1},...,p_{K}\right]$ into a stack of $L$ one-hot vector with $K$ elements, denoted as $\mathbf{P}=\left[\mathbf{I}_{p_{1}},...,\mathbf{I}_{p_{K}}\right]\in\{0,1\}^{L\times K}$. Here, $\mathbf{I}_{p_{1}}$ denotes the one-hot vector where only the $p_{1}$-th element is set to 1. As a result, the filtered event prototypes with top-$K$ scores can be summarized as $\mathbf{H}=\mathbf{P}^{\top}\mathbf{M}$. Afterwards, we resort to the perturbed maximum method [6] to construct a differentiable operator. In theory, selecting top-$K$ prototypes via subspace projection matrix $\mathbf{P}$ equals to solving a linear programming problem ,
$\mathrm{argmax}_{\mathbf{P}\in\mathcal{C}}\left\langle\mathbf{P},\mathbf{s}\mathbf{1}^{\top}\right\rangle,$ (7)
[…]
$\mathbf{P}_{\sigma}=\mathbb{E}_{P}\left[\mathrm{argmax}_{\mathbf{P}\in\mathcal{C}}\left\langle\mathbf{P},\mathbf{s}\mathbf{1}^{\top}+\sigma\mathbf{Z}\right\rangle\right],$ (8)
where $\mathbf{Z}$ is a perturbed vector sampled from the uniform Gaussian distribution and $\sigma$ serves as the hyper-parameter to control the variance of perturbation. Following [1], the Jacobian associated with Eqn. 8 can be simplified as,
$\frac{\partial\mathbf{P}_{\sigma}}{\partial\mathbf{s}}=\mathbb{E}_{P}\left[\mathrm{argmax}_{\mathbf{P}\in\mathcal{C}}\left\langle\mathbf{P},\mathbf{s}\mathbf{1}^{\top}+\sigma\mathbf{Z}\right\rangle\mathbf{Z}/\sigma\right],$ (9)
By means of Eqn. 9, the gradient from autoregressive loss in LLM would update the distribution of representation $\mathbf{H}$ and matrix $\mathbf{P}$, thereby updating our score network $\mathcal{U}(\cdot)$ via $\frac{\partial\mathbf{P}_{\sigma}}{\partial\mathbf{s}}$ according to the chain rule. As a result, our DPE module can be trained together with LLM in an end-to-end fashion, which effectively mitigates the answer-irrelevant visual nuisance in video while achieving dynamic selection in accordance with answer and question.
@@ 3.3 Compact Cooperative Encoding
We find that semantic abstraction of $\mathbf{v}_{t}$ can be achieved when spatial clustering is performed multiple times. For example, concepts such as “person” and “dog” are progressively formed from the low-level attribute or color information. We thus use multiple clustering layers to capture more abundant visual details. The output spatial object prototypes of layer $j$ is fed to layer $j+1$ for abstraction, which means that number of prototypes participating in subsequent layers reduce progressively.
Collecting all layer outputs results in our multi-grained spatial object prototypes $\mathbf{G}_{t}=\{\mathbf{Z}^{(j)}_{t}\}_{j=1}^{J}$. Here, $J$ is the total number of clustering layers.
@@
Given embeddings $\mathbf{U}_{t,b_{t}=1}$ and $\mathbf{U}_{t,b_{t}=0}$ from Cones and Rods, respectively, we combine them in a token-wise manner to obtain the dynamic embedding of the $t$-th frame,
@@ 4.1 Experimental Setup
Implementation Details. We use the pre-trained ViT-G/14 from EVA-CLIP [15] as the visual encoder to extract the features of each frame in video, and it can be further changed to other clip-based video encoders. We use pre-trained Qformer weight from InstructBLIP [12] as the textual encoder. Besides, we adopt the Vicuna-7B-1.5 model [10] as our foundational LLM. Our model is trained using 8 $\times$ NVIDIA A100 80G GPUs. See more details in the supplementary material.
@@
ST-LLM [ECCV 24][43] 7B 74.6 3.9 63.2 3.4 50.9 3.3
DynFocus (Turbo-16k) 7B 72.3 3.9 59.8 3.4 49.4 3.4
DynFocus (Turbo-0613) 7B 74.8 4.0 62.8 3.6 50.3 3.4
Table 2: Performance comparisons on VCG-Bench. $\dagger$ represents the version that first fine-tuned on all the dataset, and further post-tuning on VideoChatGPT-100K [45] with a smaller learning rate.
@@
DynFocus 7B 3.27 3.15 3.78 2.86 2.78 3.17
Table 3: Performance Comparisons on LV-Bench. Input shows the number of frames each model actually process when testing. $\dagger$ denotes the optimal results when adopting different number of $L$ on 200 input video frames.
Method Size Input ER EU KIR TG Rea Sum Overall
Short Video MLLMs
TimeChat [CVPR 24] [55] 7B 96 f 21.9 21.7 25.9 22.7 25.0 24.1 22.3
PLLaVA [73] 34B 16 f 25.0 24.9 26.2 21.4 30.0 25.9 26.1
LLaVA-NeXT [79] 34B 32 f 30.1 31.2 34.1 31.4 35.0 27.6 32.2
GPT-4o [51] - 10 f 26.5 23.7 28.3 21.4 28.0 32.8 27.0
Long Video MLLMs
MovieChat [CVPR 24] [60] 7B $\sim$10k f 21.3 23.1 25.9 22.3 24.0 17.2 22.5
LLaMA-VID [ECCV 24] [35] 13B $\sim$10k f 25.4 21.7 23.4 26.4 26.5 17.2 23.9
LWM [41] 7B $\sim$4k f 24.7 24.8 26.5 28.6 30.5 22.4 25.5
Gemini 1.5 Pro [54] - $\sim$4k f 32.1 30.9 39.3 31.8 27.0 32.8 33.1
DynFocus ($L=25,K/L=0.8$) 7B 200 f 27.9 30.3 31.2 25.4 31.8 32.8 30.4
DynFocus ($L=50,K/L=0.8$) 7B 200 f 28.6 31.8 32.6 27.2 35.3 34.4 31.8
DynFocus ($L=60,K/L=0.6$) 7B 200 f 29.9 33.7 35.1 25.5 33.3 26.2 32.6
DynFocus ($L=70,K/L=0.4$) 7B 200 f 31.8 33.5 32.6 28.7 34.8 31.3 32.9
DynFocus ($L=80,K/L=0.4$) 7B 200 f 31.1 33.5 31.6 28.6 33.8 24.1 31.8
@@
VideoChat [31] 16 f 33.0 32.0 2.31 27.0 32.1 27.6 5.01 24.3 28.6 29.2 3.66
Video-ChatGPT [ACL 24] [45] 100 f 26.9 24.0 2.31 40.3 42.0 29.9 5.48 25.1 31.1 31.3 3.90
Video-LLaMA2 [9] 16 f 54.5 41.5 2.34 39.4 33.5 35.4 5.22 18.5 25.7 35.5 3.78
VideoChat2 [CVPR 24] [32] 16 f 74.6 51.5 2.57 42.0 47.4 43.8 5.04 22.8 29.6 44.5 3.81
Video-LLaVA [36] 8 f 71.6 57.0 2.43 53.2 45.2 48.4 5.25 20.1 35.9 47.3 3.84
@@ 4.2 Evaluation on Short Video Understanding
Video-LLaVA [EMNLP 24] [37] 8 f 7B 45.3 46.1 38.0 40.7 36.2 38.1 39.9 41.6
ST-LLM [ECCV 24] [42] 16 f 7B 45.7 48.4 36.8 41.4 31.3 36.9 37.9 42.3
VideoChat2 [CVPR 24][32] 16 f 7B 48.3 52.8 37.0 39.4 33.2 39.2 39.5 43.8
Chat-UniVi  [CVPR 24] [25] - 7B 45.7 51.2 40.3 44.6 35.8 41.8 40.6 45.9
DynFocus ($L=25,K/L=0.8$) 16 f 7B 50.9 53.7 43.7 46.0 37.7 43.6 44.1 47.8
@@ 4.3 Evaluation on Long Video Understanding
To demonstrate the advantage of our dynamic cooperative setting, we conduct experiment on three newly released long-term video benchmark. The detailed description for each benchmark are elaborated in supplementary material. MLVU-Bench performance. The performance of individual task and the average performance of multi-choice task (M-Avg, within 0-100%) and generation task (G-Avg, within 0.0-10.0) are both reported in Table 4. We have following observations: (1) our DynFocus surpasses all the open-sourced video MLLMs with a clear-cut performance gain on M-Avg and G-Avg, and it nearly consistently ranks top-2 position on individual tasks. (2) For TR, AR, and VS tasks that require an thorough understanding of entire video, our method achieves the best. We attribute this to our dynamic cooperative network’s ability to balance intricate spatial details with broader temporal perception without introducing external visual nuisance. (3) Most approaches find AO and AC tasks challenging due to their sensitivity to the temporal clues, which requires recalling multiple nuanced details from lengthy videos. Although not being further fine-tuned on long-term video dataset like MovieChat, our model still performs competitively. (4) However, our method struggles with ER task that needs ego-based perspectives, likely due to the requirement for ego-centric dataset like EgoQA [13] in VideoChat2.
LV-Bench Performance. We assess six core capabilities of our model on LV-Bench: Temporal Grounding (TG), Summarization (Sum), Reasoning (Rea), Entity Recognition (ER), Event Understanding (EU), and Key Information Retrieval (KIR). The average duration of each video exceeds 1 hour. Following [65], we select several publicly evaluated methods as baselines, with results shown in Table 3. Interestingly, some methods that excel on short videos perform almost randomly in answer selection. Remarkably, our DynFocus achieves the best of 32.9% among all the open-sourced 7B models, even outperforming PLLaVA [32] with 34B parameters.
@@
VideoMME Benchmark Performance. VideoMME benchmark spans across three kinds of durations, VideoMME-S ($\sim$1.3min), VideoMME-M ($\sim$8.5min), and VideoMME-L ($\sim$0.7h). Most videos include both subtitles and audios, which helps us investigate the performance gain from additional information sources. Table 5 compares our results with other representative video MLLMs. Notably, our DynFocus consistently achieves impressive advantage across different lengths of video with subtitles and without subtitles. Specifically, it exhibits the overwhelming advantage over SOTA ST-LLM and VideoChat2. Remarkably, the version of DynFocus without subtitles reaches an overall accuracy of 44.1%, still surpassing ST-LLM with subtitles by 1.8%.